OpenStudy (anonymous):

What is the first step in solving this system of equations using elimination? y=3/4x+3

hero (hero):

One equation by itself doesn't represent a system of equations. There must be at least two equations in order for a set of equations to be called a system.

OpenStudy (anonymous):

and 2x – 4y = 6 !!

OpenStudy (anonymous):

@Hero my bad forgot to put it

hero (hero):

\(y = \frac{3}{4}x + 3\) \(2x - 4y = 6\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what is the first step

hero (hero):

Multiply both sides of the first equation by 4 to get rid of the fraction.

OpenStudy (anonymous):

write the 2nd equation in slope intercept form??

OpenStudy (anonymous):

that isn't one of the options!

hero (hero):

Well, that's the first thing I would do. But hey, different mathematicians have different methodology/philosophy on how mathematics should be done.

OpenStudy (anonymous):

true true